OpenStudy (anonymous):

The midpoint of a line segment with end points as (x, 4) and (-2, -2) is (2, 1). What is the value of x? x = 6 x = 4 x = -3 x = -2

OpenStudy (lgbasallote):

\[midpoint = \frac{x_1 + x_2}{2}\] \[2 = \frac{-2 + x}{2}\] \[4 = -2 + x\] do you see it now?

OpenStudy (anonymous):

6?

OpenStudy (anonymous):

6 - -2 = 4 so the answer would be 6 right?

OpenStudy (lgbasallote):

yup

OpenStudy (lgbasallote):

6 is right

OpenStudy (anonymous):

oh ha. that was simple. it was justy confusing to me because i thought there is suppose to only be 2 sets of numbers to do the formula

OpenStudy (lgbasallote):

yup there is...but since you're only looking for x why make it difficult :)

OpenStudy (lgbasallote):

the complete formula is \[\large midpoint = \frac{x_1 + x_2}{2} \quad , \quad \frac{y_1 + y_2}{2}\]

OpenStudy (anonymous):

oh okay that makes sense now. thankyou :D